Express vector 10 C=—a, +r cos Odo +ap in Cartesian and Cylindrical coordinates. Find C(-3, 4, 0) and C(5, 7/2, -2). Cuma
Answer #1

And Given that B = y as tor cose ao ta Now, Ba Bor art BoaO +Boad Boya 1% = 7coso Bd = 1 from spherical to cartesian :- BO CoÖ= tan-11-44-3) -- 53.130 4 5 d = cost G :) you V1-39-764)702 3 B at (-3,4, 0) . B at 15, -53.13°, 90) BE (16*3207 5(310) =( - ره رگ )8 ,35e0 حرة .,د 3 Co ۵(۰) +5(۰) و ( 62 (1) 5 - (1 + 4 + . ( دهی . (به ) .ره - رد/ری) -5 at B : 24 F4o Thank you.